OpenStudy (anonymous):

Find the value of 12(log10^2)(lne^2).

OpenStudy (anonymous):

12log^2 / log^2(10) if you mean 12(log^2base of 10)(log^2)??

OpenStudy (anonymous):

it's 12(log of 10 to the 2)(ln of e to the 2) I'm assuming the first log is just at base 10.

OpenStudy (anonymous):

log of 10 to the 2 means power of two means squared? right

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

12log^2 (10)

OpenStudy (rulnick):

It looks like angalc is on the way, but if you want a double-check, then I think what you want is: 12 (log 10^2) (ln e^2) = 12 (2 log 10) (2 ln e) = 12 (2) (2) = 48 Good luck!

OpenStudy (anonymous):

the log or the 10 is squared?? well rulnick answer is correct....kk i thought is log squared.